OpenStudy (anonymous):

Divide. (–10x3 + 30x – 20) ÷ (–5x + 5)

OpenStudy (anonymous):

@dpaInc help

OpenStudy (callisto):

For dividend, Consider f(x) = -10x^3 + 30x - 20 f(1) = -10(1)^3 + 30(1) - 20 = 0 so, (x-1) is a factor of f(x) f(x) = -10x^3 + 30x - 20 = -10 (x^3 -3x +2) By long division, x^2 + x - 2 ----------------------- x-1 ) x^3 - 3x + 2 x^3 - x^2 --------------------- x^2 - 3x x^2 - x ---------------- - 2x +2 - 2x +2 ------------ f(x) = -10x^3 + 30x - 20 = -10 (x^3 -3x +2) = -10 (x-1)(x^2 + x -2) = -10 (x-1) (x-2)(x+1) For divisor, -5x + 5 = -5 (x-1) (–10x3 + 30x – 20) ÷ (–5x + 5) = -10 (x-1) (x-2)(x+1) ÷ -5 (x-1) \[ = \frac{-10 (x-1) (x-2)(x+1)}{ -5 (x-1)}\] Can you cancel the common factors there?